The perimeter of a rectangle is 86 inches. The width of the rectangle is 5 less than two times it's length. What is the area of the rectangle? Show all work if possible


Thank you!!!

Answer by hamsanash1981@gmail.com(151)

Let length be x then, width = x - 5
perimeter = 86
2(l+b)=86
2(x+x-5)=86
4x-10 =86
4x = 96
x = 24
length = 24
Width = 24-5=> 19
area = l*b
= 24*19
=456 sq inches